OpenStudy (anonymous):

Which of these ordered pairs is a solution to the linear inequality y – 4x ≤ –6? (Points : 1) (–2, 4) (1, –2) (1, 3) (2, 3)

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@AJ01

OpenStudy (triciaal):

check each option to see if the equation is true when you use those values

OpenStudy (triciaal):

for the first one (-2, 4) x = -2 y </+ -6 + 4x y = 6 + 4(-2) = 6 - 8 = -2 -2 not = 4 therefore (-2, 4) not a solution

OpenStudy (triciaal):

next when x = 1 what is y?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

i dont know

OpenStudy (triciaal):

look at what I did replace the value of x in the equation to get the value of y then check does your value of y match the value of y given in the set? it if matches then it is a solution. if it does not match it is not a solution.

OpenStudy (triciaal):

y – 4x ≤ –6 replace x with 1

OpenStudy (anonymous):

y-4x=< -6 replace x and y a pair (x,y) then add the values 1 point (-2,4)....... 4-4(-2)=<-6 4+8=<-6 12=<-6 12 can't be less than or equal -6 so this point is not a pair of this function

OpenStudy (anonymous):

what @triciaal wrote is a nice and ease to follow

OpenStudy (triciaal):

@AJ01 thank you @ShayleeCutie do you think you can close this now?